Given that p(x) = x is a particular solution to the differential equation y'' + y = x write the generalized solution and check by verifying that the solution satisfies the equation.


1
Expert's answer
2021-09-22T00:44:53-0400

"\text{The characteristic equation for the given differential equation is given by}\\
y^2+1 = 0\\
\implies y = \pm i\\
\text{Hence the complimentary solution is given by}\\
y = C_1e^{ix} + C_2e^{-ix}\\
\implies y_c(x) = C_1\sin x + C_2 \cos x\\
\text{The general solution is given by $y_c(x) + y_p(x)$}\\
\text{$y_p(x)$ is already given, that is, x}\\
\implies y = C_1 \sin x + C_2 \cos x+ x\\
\text{Let y(x) = x be a solution, therefore $y'(x)=1$ and $y''(x) = 0$}\\
\text{Hence substituting into the given differential equation, we have}\\
y'' + y = x\\
\text{Showing that y = x is a solution of the given differential equation}"